A stone is thrown horizontally off a cliff with a velocity of 10 m/s. If it takes 5 seconds for the stone to hit the ground, how his is the cliff?
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

A stone is thrown horizontally off a cliff with a velocity of 10 m/s. If it takes 5 seconds for the stone to hit the ground, how his is the cliff?
------------------
s(t) = -9.8t^2 + vot + so
---
so is the height of the cliff.
s(t) is 0 when the stone hits the ground.
Since the stone is not thrown up and down vo is 0.
-------------------------
Find 9.8(5)^2 = 9.8*25 = 245 meters (height of the cliff)
Cheers,
Stan H.
==========================================================

Answer by Alan3354(69443)

A stone is thrown horizontally off a cliff with a velocity of 10 m/s. If it takes 5 seconds for the stone to hit the ground, how his is the cliff?
-------------------
The acceleration due to gravity is 9.8 meter/sec/sec
Velocity = acc x time, v = at
v at impact = 9.8*5 = 49 meters/sec
The average v over the 5 seconds is 24.5 m/sec
5*24.5 = 122.5 meters
-----------------
Another method:
Distance fallen = 4.9t^2 meters
s = 4.9*5^2
s = 122.5 meters
----------------
PS The horizontal velocity is irrelevant unless it is sufficiently large that the curvature of the Earth is a factor.